OpenStudy (anonymous):

cant remember how to do multiply (x+8)6x and (x+8)(x-2)

OpenStudy (anonymous):

\[a(b+c)=ab+ac\] distrubutive property of equalities

OpenStudy (mathstudent55):

The distributive property will help: (a + b)c = c(a + b) = ca + cb (x + 6)6x = 6x(x + 6) = 6x(x) + (6x)(6) Now do those two multiplications.

OpenStudy (anonymous):

so it would be 6x^2+48=84

OpenStudy (anonymous):

which is the same because \[(a+b)c=ac+bc=ca+cb=c(a+b)\]

OpenStudy (anonymous):

=]

OpenStudy (anonymous):

that only works for the first... the second is a method known as foiling \[(a\pm b)(c\pm d)=ac\pm ad\pm bc \pm bd\]

OpenStudy (mathstudent55):

6x(x + 8) = 6x(x) + (6x)(8) = 6x^2 + 48x That's it. Unlike terms can't be added together. Unlike terms are terms that have different variable parts. x^2 and x are different variable parts.

OpenStudy (anonymous):

so 27x for the swcond?

OpenStudy (mathstudent55):

Second one is the product of two binomials where you use FOIL to do it. Are you familiar with FOIL?

OpenStudy (anonymous):

if this helps you can also use the distributive \[(a\pm b)(c \pm d)=a(c\pm d)\pm b(c\pm d) \]by distributive

OpenStudy (anonymous):

which is the same as what i put above

OpenStudy (anonymous):

im not familiar to FOIL

OpenStudy (mathstudent55):